# Hierarchical Generative Adversarial Imitation Learning with Mid-level Input Generation for Autonomous Driving on Urban Environments

### v4

@@ Abstract
Deriving robust control policies for realistic urban navigation scenarios is not a trivial task. In an end-to-end approach, these policies must map high-dimensional images from the vehicle’s cameras to low-level actions such as steering and throttle. While pure Reinforcement Learning (RL) approaches are based exclusively on engineered rewards, Generative Adversarial Imitation Learning (GAIL) agents learn from expert demonstrations while interacting with the environment, which favors GAIL on tasks for which a reward signal is difficult to derive , such as autonomous driving . However, training deep networks directly from raw images on RL tasks is known to be unstable and troublesome. To deal with that, this work proposes a hierarchical GAIL-based architecture (hGAIL) which decouples representation learning from the driving task to solve the autonomous navigation of a vehicle. The proposed architecture consists of two modules: a GAN (Generative Adversarial Net) which generates an abstract mid-level input representation, which is the Bird’s-Eye View (BEV) from the surroundings of the vehicle; and the GAIL which learns to control the vehicle based on the BEV predictions from the GAN as input. hGAIL is able to learn both the policy and the mid-level representation simultaneously as the agent interacts with the environment. Our experiments made in the CARLA simulation environment have shown that GAIL exclusively from cameras (without BEV) fails to even learn the task, while hGAIL, after training exclusively on one city, was able to autonomously navigate successfully in 98% of the intersections of a new city not used in training phase.
@@ I Introduction
    extending previous work that is applicable to fixed routes [15] to a more skilled agent able to follow dynamic routes, making the vehicle able to change the route on the fly.
@@ II Related Works
II-A Behavior Cloning
The authors of works [8] and [9] utilized behavior cloning (BC) for autonomous driving (AD) in the CARLA simulator. A large dataset of human driving was collected and augmented using image processing techniques to train end-to-end policies conditioned on the desired route. Further advancements in BC were made in [26], which used a large deep ResNet network for feature extraction and fused data from camera, LiDAR, and radar to generate feature maps. The work also incorporated a probabilistic motion planner to account for uncertainties in the trajectory.
@@ II-B Reinforcement Learning (RL)
An RL agent [27] was trained and tested in a 2D simulator, CARLA, and a real-world car racing task, where the vehicle needed to repeatedly complete a predefined circular trajectory. The agent was trained using a predictive world model, which substituted the environment for a 10-state horizon. The predictive world model was trained by deploying the agent to the chosen environment, reducing the necessary environment interactions for training the agent. The project highlights the potential use of auxiliary models to address the weakness of RL in needing large amounts of environment interactions.
A different autonomous driving project [17] employed reinforcement learning (RL) with a customized reward function to train an agent based on the mid-level BEV representation as input, during the first training phase. Afterwards, using the trained agent as an online expert, they trained a second agent through apprenticeship learning, but now in an end-to-end approach with input directly from the vehicle’s cameras. The method was evaluated on the CARLA, surpassing the default benchmarks trained by behavior cloning.
@@ II-C GAIL
In [29], model-based GAIL (MGAIL) was used with a large-scale expert dataset (100,000 miles) from San Francisco city, and applied to the task of dense urban self-driving. The discriminator and generator in their MGAIL model consists of Transformer networks, which combine different types of inputs: roadgraph points, traffic light signals, and other objects’ trajectories. They introduce an hierarchical model which integrates a high-level graph-based search (to generate the intended trajectory) with a low-level transformer-based MGAIL policy. Besides, they do not learn mid-level input representations from raw camera images as hGAIL in this work, but use a pretrained independent robotic perception system to generate input features. Additionally, our controller in hGAIL learns with expert samples generated in the CARLA simulator corresponding to only 36 minutes (8 km) of driving. In theory, transfering the trained policy in hGAIL to a real-world setting would require only replacing the BEV network by another network trained on real images.
@@ III-A Conditional Generative Adversarial Networks - CGAN
Additionally, a L1 distance loss function is added to the final objective, making the generator network G also learn from the true label $y$ as it would happen in a supervised learning task, modeling low-frequency characteristics of images [22]. The resulting method, PatchGAN [22], divides the image into multiple patches to be classified by the discriminator individually and, thus, requires fewer parameters.
@@ III-B Generative Adversarial Imitation Learning - GAIL
The generator loss is also augmented with a Behavior Cloning (BC) loss ($-\mathop{\mathbb{E}}_{\pi_{E}}[\log(\pi(a|s))]$) using the available expert samples. The underlying idea is to provide the training process of the policy with useful and informative gradients, especially when the discriminator is not yet fully trained. This concept and the practical implementation follows [23].
@@ III-C Bird’s-Eye View (BEV) representation
Figure 1: a) Images from three frontal cameras located at the left, central, and right part of the vehicle, respectively. They were taken after the first few interactions of the agent in the CARLA simulation environment Each camera produces a 256x144 RGB image. b) The corresponding sparse trajectory visual input captured at the same frame. The points from the sparse trajectory and the highlighted vehicle position are plotted as circles with a radius of 10 pixels, using the same scale (pixels per meter) and perspective as the BEV representation. When the image is fed to the CGAN, it is represented with only one channel and a size of 192x192 pixels. c) The three channels of BEV image that our agent employs, computed at the same instant shown in (a). From left to right, the channels correspond to: desired route, drivable area, and lane boundaries. The last image shows all three channels combined in different colors.
@@ IV-A BEV generation with CGANs
The Conditional GAN module, used to transform the images from the frontal cameras into a top-down view representation, has two different networks named Discriminator and Generator, whose architectures can be seen in Fig. 3. In this figure, all layers from both networks are presented, and the common layers in the orange color refer to layers that exist in both networks’ architectures, even though they do not share weights (parameters).
@@ IV-A1 Input representation
IV-A2 Network Architecture
Both networks’ architectures are seen in Fig. 3, where the common layers in orange refer to layers existing in both the generator and the discriminator. Notice that they are separate networks which do not share parameters: the figure was made to not repeat equivalent layers when describing both networks.
Generator
Discriminator
@@ IV-B Policy learning with GAIL
The generator in the GAIL module iteratively seeks the $\theta$ parameters of the policy $\pi_{\theta}(.|s)$ that minimizes (2), while the discriminator seeks to maximize it. To assist the agent’s learning, loss terms for stimulating exploration are added as described after the the representations for the input, output, and architecture are presented.
@@ IV-B2 Output representation
IV-B3 Networks’ Architectures
IV-B4 Encouraring Exploration
@@
Thus, uniting $\mathcal{L}_{P}$ from policy gradient method, $\mathcal{L}_{BC}$ from Behavior Cloning Loos and (3), the total loss function for policy learning through PPO for our hGAIL agent is as follows:
$\alpha\mathcal{L}_{BC}+(1-\alpha)\mathcal{L}_{P}+\mathcal{L}_{\mathrm{ent}}+\mathcal{L}_{\mathrm{exp}}$ (7)
@@ V Experimental Results
The goal of the vehicle is to navigate autonomously in the city shown in Fig. 5 using the hGAIL agent’s architecture with mid-level BEV input generation  [^1] The code and videos for this paper are available on Github at: https://github.com/gustavokcouto/hgail.
@@ V-B Training
At every $12,288$ environment interactions (steps), the agent’s architecture is updated in a central computer: the parametrized policy using loss function (7) is trained for 20 epochs using PPO (K=20), while the GAIL’s Discriminator is trained for 2 epochs (J=2) on these $12,288$ samples; and the CGAN’s Generator for BEV with the loss function (1) is trained for 4 epochs, while its Discriminator for 4 epochs. This process corresponds to one training cycle of the full hGAIL. A new cycle will collect the next $12,288$ samples from all the actors, and execute the training as described above again. As six parallel actors are used, 2,048 steps or environment interactions per actor are recorded, totalling the 12,288 environment interactions. Thus, the episode does not have to end for a policy update to happen. It is important to note that at any given moment, any of the six actors may be interacting with the environment in different parts of the environment. Additional hyperparameters’s values can be found in Tables  I and Table II for the GAIL and GAN parts of hGAIL, respectively.
@@
TABLE I: Hyperparameters for GAIL
Description Value
Parallel environments ($N$) $6$
Initial adam step size (lr) $2.0\times 10^{-5}$
Adam step size exponential decay ($\lambda_{lr}$) $0.96$
Number of PPO epochs (K) $20$
Mini-batch size (m) $256$
Discount ($\gamma$) $0.99$
GAE parameter ($\lambda$) $0.9$
Clipping parameter ($\epsilon$) $0.2$
Value Function clipping parameter ($\epsilon_{vf}$) $0.2$
Value Function coefficient ($c_{1}$) $0.5$
Entropy coefficient ($c_{2}$) $0.01$
Exploration coefficient ($c_{3}$) $0.05$
Timesteps per epoch (T) $12288$
GAIL gamma ($\gamma_{gail}$) $0.004$
GAIL gamma decay ($\lambda_{gail}$) $1.0$
Discriminator adam step size (lr) $2.5\times 10^{-4}$
Number discriminator epochs (J) $2$
TABLE II: Hyperparameters for CGAN
Description Value
Adam step size (lr) $2.0\times 10^{-4}$
Number of GAN epochs (L) $4$
Mini-batch size (m) $32$
Patch size ($\gamma$) $(10,10)$
Resize ($\lambda$) $(192,192)$
Lambda pixel ($\epsilon$) $100$
@@ V-C Evaluation
The main evaluation environment is town02, shown in Fig. 6. It was used to test how well the hGAIL agent can generalize its driving skills to unseen, new environments. All experiments below consider agents trained exclusively in town01 environment.
Figure 6: Town02 environment of the agent, with one of the routes used to evaluate the agent trained in town02. The highlighted path has 1010 meters, 29 points in the sparse trajectory (shown as yellow dots) and 1030 points in the dense point trajectory (not shown).
@@
Figure 7: Number of committed infractions vs. environment interactions during training in town1 environment. The top (bottom) plot shows the results for the agents receiving (disregarding) the sparse trajectory as visual input (trajectory 1x192x192 in Fig.3) in the corresponding CNN. For each method (hGAIL, GAIL with real Bird’s-Eye View, GAIL from cameras), the average performance of three runs is depicted considering a stochastic policy. The shaded area represents the standard deviation. The GAIL from cameras agent fail to learn the task and keep the sum of committed infractions above zero, while the minimum of zero infractions is achieved by both hGAIL and GAIL with real BEV.
@@
The agent trained only on town1 was also evaluated at every T intersection in town2 environment, i.e., 8 different T intersections, and compared to Behavior Cloning and GAIL from cameras agents. The latter corresponds to a GAIL agent with input directly from the vehicle’s three frontal cameras instead of the mid-level BEV input. The results are summarized in Table III, whose lines presents the results for each possible turn out of 6 in total at a given T intersection (as shown in Fig. 9). Thus, each turn, covering around 100 meters, was evaluated in 8 different T intersections, totalling 48 experiments for each agent. The success percentage for each turn type is given in this table, where we can see that hGAIL can turn without failing in all intersections and for all turn types except for one top-right intersection , while BC fails 50% of the times, and GAIL from cameras fails to learn most of the required driving behavior, succeeding only in 4 turns out of 48 . This ablation of the GAN from hGAIL (which is the GAIL from cameras) shows the need for learning the mid-level input representation to succeed in this complex task. Additionally, notice that only hGAIL can finish complete trajectories (of around 1 km, as shown in Fig. 8), as both BC and GAIL from cameras fail in at least one turn type.
TABLE III: Evaluation performance for 8 T Intersections and 6 type of turns in Town2
@@
a) Refer to caption b) Refer to caption
c) Refer to caption
@@ V-D Mid-level representation learning
Figure 11: BEV generation as the agent goes through training in town2. The first three colmns shows the images from the cemeras attached to the front of the vehicle. The fourth column show the plot from the point from the route planner. The fifth column shows five BEV images computed by the simulator and are considered the target output. The following columns show the BEV images generated by the GAN from the agent’s architecture as it undergoes training, at: 12,288 environment steps (1 cycle), 208,896 environment steps ( 16 cycles), and 405,504 environment steps ( 32 cycles). One cycle is similar to the concept of epoch, and consists of the full training of hGAIL using the last 12,288 steps collected; however, each individual network of hGAIL is trained for different number of epochs in one training cycle (see Section V-B).

### v5

@@ Abstract
Deriving robust control policies for realistic urban navigation scenarios is not a trivial task. In an end-to-end approach, these policies must map high-dimensional images from the vehicle’s cameras to low-level actions such as steering and throttle. While pure Reinforcement Learning (RL) approaches are based exclusively on engineered rewards, Generative Adversarial Imitation Learning (GAIL) agents learn from expert demonstrations while interacting with the environment, which favors GAIL on tasks for which a reward signal is difficult to derive, such as autonomous driving. However, training deep networks directly from raw images on RL tasks is known to be unstable and troublesome. To deal with that, this work proposes a hierarchical GAIL-based architecture (hGAIL) which decouples representation learning from the driving task to solve the autonomous navigation of a vehicle. The proposed architecture consists of two modules: a GAN (Generative Adversarial Net) which generates an abstract mid-level input representation, which is the Bird’s-Eye View (BEV) from the surroundings of the vehicle; and the GAIL which learns to control the vehicle based on the BEV predictions from the GAN as input. hGAIL is able to learn both the policy and the mid-level representation simultaneously as the agent interacts with the environment. Our experiments made in the CARLA simulation environment have shown that GAIL exclusively from cameras (without BEV) fails to even learn the task, while hGAIL, after training exclusively on one city, was able to autonomously navigate successfully in 98% of the intersections of a new city not used in training phase. Videos and code available at: https://sites.google.com/view/hgail.
@@ I Introduction
    extending previous work that is applicable to fixed routes [15] to a more skilled agent able to follow dynamic routes, making the vehicle able to change the route on the fly. In addition, this work also tackles navigation in cities with pedestrians and other vehicles, and respecting traffic lights, unlike [15].
@@ II Related Works
II-A Behavior Cloning (BC)
The authors of works [8] and [9] utilized Behavior Cloning (BC) for Autonomous Driving (AD) in the CARLA simulator. A large dataset of human driving was collected and augmented using image processing techniques to train end-to-end policies conditioned on the desired route. Further advancements in BC were made in [26], which used a large deep ResNet network for feature extraction and fused data from camera, LiDAR, and radar to generate feature maps.
@@ II-B Reinforcement Learning (RL)
In [27], an RL agent was trained and tested in a 2D simulator, CARLA, and a real-world car racing task, highlighting the potential use of auxiliary models to address the weakness of RL in needing large amounts of environment interactions. In [17], RL is employed with a customized reward function to train an agent based on the mid-level BEV representation as input, during the first training phase. Afterwards, using the trained agent as an online expert, they trained a second agent through apprenticeship learning, but now in an end-to-end approach with input directly from the vehicle’s cameras. Their method evaluated using CARLA surpassed the default benchmarks trained by BC.
@@ II-C GAIL
In [29], model-based GAIL (MGAIL) was used with a large-scale expert dataset (100,000 miles) from San Francisco city, and applied to the task of dense urban self-driving. The discriminator and generator in their MGAIL model consists of Transformer networks, which combine different types of inputs: roadgraph points, traffic light signals, and other objects’ trajectories. They introduce an hierarchical model which integrates a high-level graph-based search (to generate the intended trajectory) with a low-level transformer-based MGAIL policy. Besides, they do not learn mid-level input representations from raw camera images as hGAIL in this work, but use a pretrained independent robotic perception system to generate input features. Additionally, our controller in hGAIL learns with expert samples generated in the CARLA simulator corresponding to only 36 minutes (8 km) of driving. In theory, transferring the trained policy in hGAIL to a real-world setting would require only replacing the BEV network by another network trained on real images.
@@ III-A Conditional Generative Adversarial Networks - CGAN
Additionally, a L1 distance loss function is added to the final objective, making the generator network G also learn from the true label $y$ as it would happen in a supervised learning task, modeling low-frequency characteristics of images. The resulting method, PatchGAN [22], divides the image into multiple patches to be classified by the discriminator individually and, thus, requires fewer parameters.
@@ III-B Generative Adversarial Imitation Learning - GAIL
Mathematically, GAIL 111 More details on GAIL are given in Appendix VI-A. finds a saddle point $(\pi,D)$ of the expression:
@@
where $D:S\times A\rightarrow(0,1)$, $S$ is the state space, $A$ is the action space; $\pi_{E}$ is the expert policy; $\pi$ is the agent’s policy ; $H(\pi)$ is a policy regularizer controlled by $\lambda>=0$ [31]. The discriminator D will try to increase (2), while $\pi$ seeks to minimize it; and $L_{gp}$ is a loss that penalizes the gradient’s norm, constraining the discriminator network to the 1-Lipschitz function space, according to [32]. The above equation is the Wasserstein loss, used to alleviate vanishing gradient and mode collapse problems, using the Wasserstein distance [32] between the policy distribution and expert distribution, as also done in [33, 34].
@@ III-C Bird’s-Eye View (BEV) representation
Figure 1: a) Images from three frontal cameras located at the left, central, and right part of the vehicle, respectively. They were taken after the first few interactions of the agent in the CARLA simulation environment. Each camera produces a 256x144 RGB image. b) The corresponding sparse trajectory visual input captured at the same frame. The points from the sparse trajectory and the highlighted vehicle position are plotted as circles with a radius of 10 pixels, using the same scale (pixels per meter) and perspective as the BEV representation. When the image is fed to the CGAN, it is represented with only one channel and a size of 192x192 pixels. c) The three channels of BEV image that our agent employs, computed at the same instant shown in (a). From left to right, the channels correspond to: desired route, drivable area, and lane boundaries. The last image shows all three channels combined in different colors.
@@ IV-A BEV generation with CGANs
The Conditional GAN module, used to transform the images from the frontal cameras into a top-down view representation, has two different networks named Discriminator and Generator, whose architectures can be seen in Appendix VI-B .
@@ IV-A1 Input representation
The architectures of both Generator and Discriminator are described in Appendix VI-B.
@@ IV-B Policy learning with GAIL
The generator in the GAIL module iteratively seeks the $\theta$ parameters of the policy $\pi_{\theta}(.|s)$ that minimizes (2), while the discriminator seeks to maximize it. To assist the agent’s learning, loss terms for stimulating exploration are added as described after the representations for the input, output, and architecture are presented.
@@ IV-B2 Output representation
The architectures of both Generator and Discriminator are described in Appendix VI-C.
IV-B3 BC augmentation
The generator loss in GAIL is also augmented with a Behavior Cloning (BC) loss $\mathcal{L}_{BC}$ ($-\mathop{\mathbb{E}}_{\pi_{E}}[\log(\pi(a|s))]$) using the available expert samples. The underlying idea is to provide the training process of the policy with useful and informative gradients, especially when the discriminator is not yet fully trained. This concept and the practical implementation follows [23].
IV-B4 Encouraging Exploration
@@
Thus, uniting the Proximal Policy Optimization (PPO) loss $\mathcal{L}_{P}$ [36] that minimizes (2), the Behavior Cloning loss $\mathcal{L}_{BC}$ [23] and the entropy and exploration terms in (3), the total loss function for policy learning of all GAIL-based agents in this work is as follows:
$\alpha\mathcal{L}_{BC}+(1-\alpha)\mathcal{L}_{P}+\mathcal{L}_{\mathrm{ent}}+\mathcal{L}_{\mathrm{exp}},$ (7)
where $\alpha$ controls the weight of the BC term $\mathcal{L}_{BC}$ with respect to the PPO loss $\mathcal{L}_{P}$ during training.
@@ V Experimental Results
The goal of the vehicle is to learn to navigate autonomously in the city shown in Fig. 3 using the hGAIL agent’s architecture with mid-level BEV input generation, and subsequently generalize its driving skill to a new town.
@@ V-B Training
At every $12,288$ environment interactions (steps), the agent’s architecture is updated in a central computer: the parametrized policy using loss function (7) is trained for 20 epochs using PPO (K=20), while the GAIL’s Discriminator is trained for 2 epochs (J=2) on these $12,288$ samples; and the CGAN’s Generator for BEV with the loss function (1) is trained for 4 epochs, while its Discriminator for 4 epochs. This process corresponds to one training cycle of the full hGAIL. A new cycle will collect the next $12,288$ samples from all the actors, and execute the training as described above again. As six parallel actors are used, 2,048 steps or environment interactions per actor are recorded, totalling the 12,288 environment interactions. Thus, the episode does not have to end for a policy update to happen. It is important to note that at any given moment, any of the six actors may be interacting with the environment in different parts of the environment. Additional hyperparameters’s values can be found in Appendix VI-D.
@@ V-C Evaluation
The main evaluation environment is town2, shown in Appendix VI-F . It was used to test how well the hGAIL agent can generalize its driving skills to unseen, new environments. All experiments below consider agents trained exclusively in town01 environment.
@@
Figure 4: Number of committed infractions vs. environment interactions during training in town1 environment. The top (bottom) plot shows the results for the agents receiving (disregarding) the sparse trajectory as visual input (trajectory 1x192x192 in Fig.10) in the corresponding CNN. For each method (hGAIL, GAIL with real Bird’s-Eye View, GAIL from cameras), the average performance of three runs is depicted considering a stochastic policy. The shaded area represents the standard deviation. The GAIL from cameras agent fails to learn the task and keep the sum of committed infractions close to zero, while the goal of zero infractions is achieved by both hGAIL and GAIL with real BEV.
@@
The agent trained only on town1 was also evaluated at every T intersection in town2 environment, i.e., 8 different T intersections, and compared to Behavior Cloning and GAIL from cameras agents. The latter corresponds to a GAIL agent with input directly from the vehicle’s three frontal cameras instead of the mid-level BEV input. The results are summarized in Table I, whose lines presents the results for each possible turn out of 6 in total at a given T intersection (as shown in Fig. 6). Thus, each turn,covering around 100 meters, was evaluated in 8 different T intersections, totalling 48 experiments for each agent. The success percentage for each turn type is given in this table, where we can see that hGAIL can turn without failing in all intersections and for all turn types except for one top-right intersection, while BC fails 50% of the times, and GAIL from cameras fails to learn most of the required driving behavior, succeeding only in 4 turns out of 48. This ablation of the GAN from hGAIL (which is the GAIL from cameras) shows the need for learning the mid-level input representation to succeed in this complex task. Additionally, notice that only hGAIL can finish complete trajectories (of around 1 km, as shown in Fig. 5), as both BC and GAIL from cameras fail in at least one turn type.
TABLE I: Evaluation performance for 8 T Intersections and 6 type of turns in Town2
@@ V-D Mid-level representation learning
Figure 7: BEV generation in town2 as the agent goes through training in town1 . The first three columns shows the images from the cameras attached to the front of the vehicle. The fourth column shows the sparse trajectory from the route planner. The fifth column shows five BEV images computed by the simulator and are considered the target output. The following columns show the BEV images generated by the GAN from the agent’s architecture as it undergoes training, at: 12,288 environment steps (1 cycle), 122,880 environment steps (10 cycles), 245,760 environment steps (20 cycles), 614,400 environment steps (50 cycles), and 1,105,920 environment steps (90 cycles). One cycle is similar to the concept of epoch, and consists of the full training of hGAIL using the last 12,288 steps collected; see Section V-B for more details.
V-E Dynamic environment
Here, we extend the hGAIL agent’s architecture to allow for environments with traffic lights, pedestrians and other vehicles. Initially, to make the agent able to learn to respect traffic lights, it was necessary to employ importance weighting in the BC term ($\mathcal{L}_{BC}$) of the loss function in (7) so that each sample in the expert set is weighted according to its importance. These weights were found by computing the inverse of the output of a Kernel Density Estimator (KDE) applied on the target outputs (action values), so that less frequent actions (e.g., accelerating as the light turns green) are given more weight in the BC loss. The KDE used a Gaussian Kernel with bandwidth equal to 0.2. Additionally, three extra binary inputs were used for the GAIL policy in hGAIL, denoting the detection of traffic lights, pedestrians and vehicles in front of the cameras. Also, the weight of the BC term in (7) was set to $\alpha=0.012$, and the number of discriminator epochs ($J$) to 4.
The expert dataset was built similarly to the previous experiment, with the same trajectories, as described in Section V-A, except that now the expert respects traffic lights and breaks when pedestrians and other vehicles appear close enough ahead of its vehicle, which increases the dataset size. We trained two agents from scratch for each method, i.e., hGAIL, GAIL with real BEV, and GAIL from camera. The last two methods also employed importance weighting and three extra inputs as done in hGAIL, but processed inputs directly from the real BEV of the scene for the second method, and from the three frontal cameras for the last method. The training evolution can be seen in Fig. 8, where the first two agents converge to a low infraction rate in town01. The evaluation results can be seen in Table II for both training town1 and test town2 environments, showing the mean and the standard deviations over all routes in each town and the two agents trained for each method . Even though the GAIL agent has a slightly better Route Completion, its Driving Score (22%) is worse than the score (45%) of the hGAIL agent in town02 because the former commits more infractions and collisions. It could be that the noisy predicted BEV that hGAIL uses as input to the policy makes it more robust when learning to avoid infractions in comparison to the GAIL with real BEV agent. We can also notice that the expert fails sometimes. Particularly, hGAIL surpass the expert for the Red Light Infraction in town01, and it gets similar performance to the expert in other metrics. The GAIL agent trained with input directly from camera presented the worst performance, with 28% and 5% average driving scores for town01 and town02, respectively.
Figure 8: Number of infractions for hGAIL, GAIL from real BEV, and GAIL from camera agents during training in town1 with pedestrians, vehicles and traffic lights.
TABLE II: Driving performance and infraction analysis for different agents on Town01 and Town02.
Agent Town Metrics
Driving Route Infraction Collisions Collisions Red Light Agent
Score Completion Penalty Vehicle Pedestrian Infraction Blocked
(%, $\uparrow$)a (%, $\uparrow$) (%, $\uparrow$) (#/Km, $\downarrow$) (#/Km, $\downarrow$) (#/Km, $\downarrow$) (#/Km, $\downarrow$)
Expert town01 86 $\pm$ 9 98 $\pm$ 2 87 $\pm$ 8 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.62 $\pm$ 0.43 0.21 $\pm$ 0.21
town02 46 $\pm$ 16 82 $\pm$ 1 60 $\pm$ 18 0.31 $\pm$ 0.14 0.0 $\pm$ 0.0 2.03 $\pm$ 1.32 1.17 $\pm$ 0.29
hGAIL town01 82 $\pm$ 1 95 $\pm$ 6 86 $\pm$ 4 0.14 $\pm$ 0.14 0.09 $\pm$ 0.09 0.48 $\pm$ 0.11 0.26 $\pm$ 0.26
town02 45 $\pm$ 2 66 $\pm$ 8 65 $\pm$ 3 0.98 $\pm$ 0.49 0.17 $\pm$ 0.17 2.24 $\pm$ 0.95 2.00 $\pm$ 0.51
GAIL town01 67 $\pm$ 1 97 $\pm$ 0 70 $\pm$ 1 0.76 $\pm$ 0.27 0.05 $\pm$ 0.05 0.53 $\pm$ 0.09 0.13 $\pm$ 0.0
town02 22 $\pm$ 2 72 $\pm$ 1 31 $\pm$ 5 3.93 $\pm$ 1.13 0.08 $\pm$ 0.08 2.24 $\pm$ 0.93 1.62 $\pm$ 0.13
GAIL town01 28 $\pm$ 1 50 $\pm$ 6 61 $\pm$ 10 0.36 $\pm$ 0.0 0.35 $\pm$ 0.35 2.63 $\pm$ 0.46 8.32 $\pm$ 4.02
from cam town02 5 $\pm$ 2 10 $\pm$ 3 48 $\pm$ 4 10.76 $\pm$ 6.80 0.62 $\pm$ 0.62 0.62 $\pm$ 0.62 18.00 $\pm$ 3.75
  • a
    The Driving Score, a metric from CARLA Leaderboard, is calculated by the product of Route Completion (the percentage of the route distance completed) and an Infraction Penalty which accounts for all violations committed. For instance, if an agent ran a red light and collided with a pedestrian during a route and the penalty for running one red light was 0.7 and for colliding with a pedestrian is 0.5, the penalty would be $0.7\times 0.5=0.35$.
In Fig. 9, we can observe that the hGAIL agent can generate the three BEV channels for navigation in spite of the pedestrians and vehicles that may pass in front of the cameras. Thus, it learns to predict what is the road ahead in the BEV representation as if the vehicle or pedestrian would not be in front of the camera blocking the passage and the view.
Figure 9: BEV generation in a scene with two cars captured by the vehicle’s camera. Left: central camera image input to hGAIL. Right: predicted BEV output by the CGAN in hGAIL.